# Show Me the World in My Language: Establishing the First Baseline for Scene-Text to Scene-Text Translation

### v2

@@ 6 Experiments
Table 2: User Study to assess visual translation. We report mean translation quality (TQ), Readability (R), Perception Quality (PQ), and Source Style Preservation (SSP). The user study is based on evaluations by four users who are proficient in Hindi and English. The users evaluated the output on a four-point Likert scale, with four being the best quality output. D.E: Design Enhancements. Please refer Section 5.2 for more details.
Figure 6: A selection of visual translation results for proposed baseline variants for Eng $\rightarrow$ Hin in row: 1–6 and Hin $\rightarrow$ Eng in row 7–8. Here, we show (left to right) Original Image, results of baseline B-1, B-2, B-6, B-7 (please refer to Section 3.2 for details about these baseline variants). We observe that, B-7 which uses SRnet++ for scene text synthesis and proposed design enhancements is clearly superior in visual translation. Native Hindi speakers can observe that IndicTrans2 (used in B-6 and B-7) produces superior translations, and the design enhancements in B-7 result in grammatically correct translations.
@@ 6.1 Qualitative Results
We show a selection of results for variants of the proposed baseline in Figure 6.
The results indicate the merits/demerits of various choices. Usage of IndicTrans2 as the translator instead of M2M improves translation to a large extent, also enables transliterating words when necessary. By using Controlnet-based model for erasing all text regions in the image in SRNet++ instead of precariously erasing text from word crops as done by Mostel or Srnet, we ensure complete erasing of source text. The rendering of target text is also more clear and with less patchy effect. Design enhancements, particularly translating at paragraph level instead of word level improve the translation correctness by taking care of language-specific ordering of words.
@@ 6.2 Limitations
The limitations of the proposed baselines are as follows: (i) It has limited success in visually translating curved or occluded Hindi texts, partly because unlike English, scene text detection and recognition for Indian languages are still in their infancy. (ii) Another limitation is the trade-off between image quality and translation quality. If we use the design enhancements, we translate entire sentences and essentially approximate the position and size of each word in the translated sentence before rendering. These approximations lead to slight blurring of certain words as shown in Figure 6 which leads to drop in perceptual quality. However, we tolerate these issues as translation quality is preferred over image quality as long as the text is readable. (iii) Proper alignment of generated text so that they appear natural remains a challenge. Therefore, the baselines have limited success for translating longer sentences/phrases. (iv) Our current baselines do not leverage visual cues from the scene to perform translation. As a result, they often fall short in distinguishing between transliteration and translation, especially for brand names. We leave addressing these limitations as future scope of this work.
@@ 7 Conclusion
We have presented a comprehensive study for the task of visual translation. We proposed baselines for visual translation utilizing state-of-the-art approaches and their enhancements across different modules. The baselines show promising results for visually translating English-to-Hindi as well as Hindi-to-English scene text images. Nevertheless, It is evident that visual translation presents numerous challenges, and addressing all of them is beyond the scope of a single paper. We hope that introducing this task, along with the dataset, baseline, and performance metrics, will inspire the research community to develop advanced models for visual translation.
Acknowledgement: This work was partly supported by MeitY, Government of India (Project Number: S/MeitY/AM/20210114) under NLTM-Bhashini.

### v3

@@ 6 Experiments
Table 2: User Study to evaluate baselines for visual translation. We report mean Translation Quality (TQ), Readability (R), Perception Quality (PQ), and Source Style Preservation (SSP). Four fluent Hindi-English speakers rated the output on a four-point Likert scale, with 4 being the highest quality. D.E.: Design Enhancements. For more details please refer to Section 5.2.
Figure 6: A selection of visual translation results for proposed baseline variants for Eng $\rightarrow$ Hin in row 1–6 and Hin $\rightarrow$ Eng in row 7–8. Here, we show (left to right) the original image and results of best-performing baselines B-1, B-2, B-6, and B-7 (please refer to Section 3.2 for details about these baseline variants). We observe that B-7, which uses SRnet++ for scene text synthesis and proposed design enhancements, is clearly superior in visual translation. Native Hindi speakers can find that IndicTrans2 (used in B-6 and B-7) produces superior translations, and the design enhancements in B-7 result in grammatically correct translations.
@@ 6.1 Qualitative Results
We show a selection of visual results for the proposed baseline variants in Fig. 6. The illustrated results indicate the merits/demerits of various choices. The use of IndicTrans2 as the translator instead of M2M improves translation to a large extent and also enables the transliteration of words when necessary. By using a ControlNet-based model for erasing scene-text regions in the image, in SRNet++ instead of precariously erasing text from word crops as done by MOSTEL or SRNet, we ensure complete erasing of the source text. The rendering of the target text is also clearer and has a less patchy effect. Design enhancements, particularly translating at paragraph level instead of word level, improve the translation correctness by taking care of language-specific ordering of words.
Figure 7: Comparison of our proposed baseline (B-7) with Google Translate for images (a commercial application). For more details please refer to Section 6.2.
6.2 Comparison with Commercial Systems
Google Translate for images [^4] https://translate.google.com/?op=images is a commercial system that also handles scene text-to-scene text translation. However, it is closed-source and only available through a web interface, with no free API support. As a result, we do not include it in our quantitative comparison. Nevertheless, Fig. 7 presents some qualitative comparisons, illustrating that even Google Translate is not without flaws.
@@ 6.3 Limitations
The limitations of the proposed baselines are as follows: (i) It has limited success in visually translating curved or occluded Hindi texts, partly because, unlike English, scene text detection and recognition for Indian languages are still in their infancy. (ii) There is a trade-off between image and translation quality. Design enhancements allow for sentence-level translation, but approximations in word positioning and size can cause slight blurring, as illustrated in Fig. 6. While these issues affect perceptual quality, we prioritize translation accuracy over image sharpness as long as the text remains readable. (iii) Ensuring the natural alignment of generated text in the scene is challenging. Therefore, the baselines have limited success in translating longer sentences or phrases. (iv) Our baselines do not utilize visual cues from the scene, which impairs their ability to choose between transliteration and translation, particularly for brand names. Additionally, the absence of an automatic metric for evaluating source style preservation or the visual consistency between source and generated scene text, such as font, orientation, and style, limits our current evaluation framework. Addressing these limitations is an important direction for future research.
@@ 7 Conclusion
We have presented a comprehensive study for the task of visual translation by proposing a series of baselines that utilize state-of-the-art approaches and their enhancements across various modules. Our baselines demonstrate promising results for translating scene text images between English and Hindi. However, it is evident that visual translation remains challenging, and addressing all of its complexities extends beyond the scope of this single paper. We hope that introducing this task, along with the dataset, baseline, and performance metrics, will inspire the research community to develop advanced models for visual translation.
Acknowledgement
This work was partly supported by MeitY, Government of India under NLTM-Bhashini.